It takes 7.3 mJ to move a charge of 4.7 μC from Point A to Point C in an electric field. What is the potential difference between the two points?
1
Expert's answer
2020-03-02T09:12:01-0500

The work done by an electric field


"W=qV."

Hence, the potential difference


"V=\frac{W}{q}=\frac{7.3\times 10^{-3}}{4.7\times 10^{-6}}=1553\:\rm V."